OpenStudy (butterflydreamer):

Consider the geometric series sin(2x) + sin(2x)cos(2x) + sin(2x)cos^2(2x) for 0

OpenStudy (turingtest):

do you know the formuala for a geometric sum?

OpenStudy (butterflydreamer):

Is it Sn = a(1-r^n) /1-r ?

OpenStudy (turingtest):

for an infinite series it is\[\sum_0^\infty ar^n=\frac a{1-r},~if~|r|<1\]

OpenStudy (turingtest):

what is \(r\) in your case? is \(|r|<1\) ?

OpenStudy (butterflydreamer):

r would be cos2x .. i think ?

OpenStudy (turingtest):

yes, and is that <1 on the interval (0,pi/2) ?

OpenStudy (butterflydreamer):

-1<r<1 ?

OpenStudy (turingtest):

that is the requirement to be able to use the formula I gave above is -1<cos(2x)<1 in the interval 0<x<pi/2 ?

OpenStudy (butterflydreamer):

yes!

OpenStudy (turingtest):

awesome! then we can just plug into our formula for \(a\) and \(r\)

OpenStudy (butterflydreamer):

ohhh okay! I think i understand it now :) Thank you :D !!

OpenStudy (turingtest):

you're welcome :)